OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 2) = 12

OpenStudy (anonymous):

\[4(x-1)-(x+2)=12\]\[4x-4-(x+2)=12\]\[4x-4-x-2=12\]\[3x-6=12\]\[3x=18\to x=6\]

OpenStudy (anonymous):

Aw thank you!!